Let A = \(\\mathrm{\\{}\)1,2\(\\mathrm{\\}}\), B = \(\\mathrm{\\{}\)a,b,c\(\\mathrm{\\}}\), C = \(\\mathrm{\\{}\)c,d\(\\mathrm{\\}}\). Find \(A\\ \\times\\ (B\\cap\\ C). \)
a.\\(\\left[\\begin{array}{cc} 1 & 0\\\\ 0 & 1 \\end{array}\\right]\\)
b.\\(\\left[\\begin{array}{cc} 1 & 1\\\\ 1 & 1 \\end{array}\\right]\\)
c.\\(\\left[\\begin{array}{cc} -1 & 1\\\\ 1 & -1 \\end{array}\\right]\\)
d.\\(\\left[\\begin{array}{cc} 0 & 1\\\\ 1 & 0 \\end{array}\\right]\\)
1
Expert's answer
2020-05-25T20:24:47-0400

"B=(a,b,c)\\
C=(c,d))\\
B\cap C=(c)\\
A=(1,2)\\
A\times (c)=(1,2)\times(c)"

a) "c=\begin{pmatrix}
 1 & 0 \\
 0 & 1
\end{pmatrix}"

"A\times c=\begin{pmatrix}
 1 &2
\end{pmatrix}\cdot\begin{pmatrix}
 1 &0 \\
 0 & 1
\end{pmatrix}=\begin{pmatrix}
 1 & 2 
\end{pmatrix}"

b)

"c=\begin{pmatrix}
 1 & 1 \\
 1 & 1
\end{pmatrix}\\
A\times c=\begin{pmatrix}
 1 &2
\end{pmatrix}\cdot\begin{pmatrix}
 1 &1 \\
 1 & 1
\end{pmatrix}=\begin{pmatrix}
 3 & 3 
\end{pmatrix}"

c)

"c=\begin{pmatrix}
 -1 & 1 \\
 1 & -1
\end{pmatrix}\\
A\times c=\begin{pmatrix}
 1 &2
\end{pmatrix}\cdot\begin{pmatrix}
 -1 &1 \\
 1 & -1
\end{pmatrix}=\begin{pmatrix}
 1 & -1 
\end{pmatrix}"

d)

"c=\begin{pmatrix}
 0 & 1 \\
 1 & 0
\end{pmatrix}\\
A\times c=\begin{pmatrix}
 1 &2
\end{pmatrix}\cdot\begin{pmatrix}
 0 &1 \\
 1 & 0
\end{pmatrix}=\begin{pmatrix}
 2 & 1 
\end{pmatrix}"